OpenStudy (pokemon23):

The base of an isosceles triangle exceeds each of the equal sides by 3 meters. If the perimeter is 27 meters, find each side of the triangle.

OpenStudy (pokemon23):

seriously I suck at math?

OpenStudy (pokemon23):

how?

OpenStudy (anonymous):

OK, so, the iso. triangle means the two non-base sides are the same length. so b + b+3 + b+3 = 27

OpenStudy (anonymous):

the first b is the base, the second b+3 is the second side and the other b+3 is the other side; you know they must all sum to 27

OpenStudy (pokemon23):

ok

OpenStudy (mr.math):

The perimeter of a triangle is the its sides x, x+3 and x+3, \[x+x+3+x+3=27 \implies 3x+6=27 \implies 3x=21 \implies x=7.\]

OpenStudy (pokemon23):

fancy

OpenStudy (mr.math):

So the sides are \(7, 10, 10\).

OpenStudy (anonymous):

solution: (x+3) + x + X = 27 3x+3 = 27 x + 8 thus: the sides are 8, 8, 11

OpenStudy (pokemon23):

ok bye dude gtg to bed

OpenStudy (anonymous):

it's not 7, 7, 10, considering that the given states that the base side is longer than the equal sides. . .

OpenStudy (pokemon23):

dcrose is it 7,10,10

OpenStudy (pokemon23):

or 8,8,11

OpenStudy (anonymous):

the base is bigger so its 8,8,11

OpenStudy (pokemon23):

both equal =27

OpenStudy (anonymous):

another medal.....????

OpenStudy (anonymous):

oh yeah, my bad, i read it as the base being shorter too

OpenStudy (pokemon23):

show how do you get 8?

OpenStudy (pokemon23):

its say by 3meters

OpenStudy (pokemon23):

i saw how you got 7 but I don't understand how you 8?

OpenStudy (anonymous):

the base is the longer one, right? so how long are the sides, in terms of the base B?

OpenStudy (pokemon23):

add 3

OpenStudy (anonymous):

x+x+(X+3)=27 3x+3 =27 3x =27 X=8 thus: 8, 8, 11 i fail also at first.

OpenStudy (anonymous):

the base is longer, not shorter

OpenStudy (pokemon23):

I don't get it how do you get 8?

OpenStudy (pokemon23):

brb I'm going to log back

OpenStudy (anonymous):

yeah, the base is longer that's y we got 11. given isosceles triangle where two sides are equal, and that's 8 and 8..

OpenStudy (pokemon23):

type up the answer I got to go sleep (robotobey) i will be back in the morning ;) thanks for everything special thanks to (robotobey,dcrose, Mr.math, ktwlown,jerwyn gayo)

OpenStudy (anonymous):

The base of an isosceles triangle exceeds each of the equal sides by 3 meters. If the perimeter is 27 meters, find each side of the triangle. ---------------------------------------------- Let the base length of the isosceles triangle be "b" and the length of each of the other two sides be "s". An isosceles triangle is a triangle where two of the sides are equal in length. The word base is used here to refer to the third side. The perimeter, "p", is the sum of the three side lengths. p = b +2s From the problem statement, p = 27. 27 = b + 2s From the problem statement, the base, b, is longer than the length of either of the other tow sides by 3. So, b = s + 3 Replace the b in the equation, 27 = b + 2s, with (s+3). 27 = (s+3) + 2s 27 = s+3+2s 27 = 3s+3 24 = 3s 8 = s s = 8 We know from 8 lines back that b = s + 3. Replace s with 8. b = s + 3 b = 8 + 3 b = 11 Do the solutions check out with the problem statement? p = b + 2 s 27 = 11 + 2*8 ? 27 = 11 + 16 ? 27 = 27 ? Yes.

OpenStudy (anonymous):

yeah. . . correct.

OpenStudy (preetha):

Good work. The take home is that you have to read the question really carefully. Thanks for the careful thinking.